OpenStudy (anonymous):

The enrollment at a college went up from 625 to 710. what is the percent of increase?

OpenStudy (jdoe0001):

well, how much it went up anyway?

OpenStudy (tennis5518):

710-625 = 85 divide 85 by 625 85/625 = .136 13.6%

OpenStudy (tennis5518):

hope that help

OpenStudy (anonymous):

Formula for percent increase:\[ \frac{x_2-x_1}{x_1} \]

OpenStudy (tennis5518):

it is right i did this! :E